OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b?

OpenStudy (radar):

Welcome to OC. You have asked an algebra question. The question has already labeled the two numbers: One number is called "a", the other is called b.. the next thing given is "the difference of two numbers, a and b is 21. Using algebra we then can write a - b = 21. We are further given expressed in algebra: 5a - 2b =18. The key to this is a and b have the same value in both expressions. You have two equations and 2 unknowns. It is a system: a - b = 21 and 5a - 2b = 18. Use the substitution method to solve. in first equation express a in terms of b like: a = b + 21, now substitute for a in the 2nd equation: 5(b + 21) - 2b = 18 solve for b 5b + 105 -2b = 18 3b = 18 - 105 3b = -87 b=-29 Now substitute the value of (-29) of b in the first equation. a- (-29) = 21 a + 29 = 21 a= -8 As always you want to verify: -8 - (-29) = 21, getting -8 + 29 = 21 21=21 First one checks out Now do 2nd one 5(-8) - (2*-29) = 18 -40 - (-58) = 18 -40 +58 = 18 18=18 Also checks.